2. At one point in a pipeline the water’s speed is 3.00 m/s and the gauge...

2.

At one point in a pipeline the water’s speed is 3.00 m/s and the gauge pressure is 5.00∙104 Pa. Find the gauge pressure at a second point in the line, 11.0 m lower than the first, if the pipe diameter at the second point is twice that at the first.

Note: density of water is 1000 kg/m3.
Answer #1

4 A V = A. di v = TV2 de du ve d = (22, 32 ₂ = 4 - 3 wols P, + 1 2 + 18 h, = P + 2 2 2 + 12 h 5x104+ £ X1000X (334+0 = P2+ x1